# BOASF: A Unified Framework for Speeding up Automatic Machine Learning via Adaptive Successive Filtering

### v1

@@ 2.1.2 Hyperparameter Optimization
Given a learning algorithm $A$, the hyperparameters of $A$ also plays an important role in the final prediction performance. Suppose that algorithm $A$ contains $n$ hyperparameters $\lambda_{1},\cdots,\lambda_{n}$ with domains $\Lambda_{1},\cdots,\Lambda_{n}$, the hyperparameter space $\Lambda$ is the cross-product of these domains: $\Lambda=\Lambda_{1}\times\cdots\times\Lambda_{n}$. Given such a hyperparameter space $\Lambda$, we define the hyperparameter optimization problem as follows:
@@ 3.2 Adaptive Arm Filtering
Gaussian UCB takes the statistical data of historical results into consideration. Note that each arm is evaluated multiple times in a single evaluation round. Let $N$ denotes the number of tries. The formal definition of Gaussian UCB can be expressed as follows:
@@ 3.3 Adaptive Resource Allocation
[…]
@@ 5.2 Model Selection
Dataset $R$=2 hours, $r$=3 $R$=4 hours, $r$=3
@@ 5.3 Hyperparameter Optimization
BOHB is the state-of-the-art hyperparameter optimization method that outperforms random search [4] and HyperBand [19]. We compared these three methods on two widely-used models: random forest and logistic regression. Similar to model selection, different time limits (222 h and 444 h) are used to evaluate hyperparameter optimization. The hyperparameters in BOASF and BOHB are $c=2,r=3$ and $b_{min}=\frac{1}{27},b_{max}=1,\eta=3$ respectively. The random forest model is configured with 5 hyperparameters:
@@
  • max features, uniform float in [0.5,1][0.5,1][ 0.5 , 1 ].
  • min samples split, uniform integer in [2,21][2,21][ 2 , 21 ].
  • min samples leaf, uniform integer in [1,21][1,21][ 1 , 21 ].
@@
  • C, uniform float in $[1\mathrm{e}{-4},1\mathrm{e}{4}]$.
  • max iteration, uniform integer in [50,500][50,500][ 50 , 500 ].
We divided the range of each hyperparameter into 222 intervals. Therefore, BOASF has $2^{5}$ and $2^{3}$ arms respectively. Table 2 and Table 3 shows that BOASF achieved better balanced accuracy score than Hyperopt-full and BOHB on both the logistic regression and random forest settings statistically. The experimental results between BOASF and Hyperopt-full demonstrate that BOASF along with the sub-space partitioning strategy is effective for hyperparameter optimization.
@@ 5.4 Internal Framework Study
Dataset 222 hours 444 hours
@@
Dataset 222 hours 444 hours
@@ 7 DISCUSSION AND CONCLUSION
There are several improvements to do in our method in the future. Firstly, the UCB score works for a random variable range from 0 to 111. Those metrics which do not meet this condition need to be readjusted to this range. Secondly, we use MinMaxScaler to transfer the UCB score to filtering probability since it guarantees a property that the lowest score will be discarded and the highest score will advance to the next round. Other rescaling methods which has this property still work in our method. Thirdly and finally, th TPE method is used to tune the hyperparameters of each arm. We can compare it with other Bayesian Optimizaiton methods.

### v2

@@ 2.1.2 Hyperparameter Optimization
Given a learning algorithm AA, the hyperparameters of AA also plays an important role in the final prediction performance. Suppose that algorithm AA contains $n$ hyperparameters $\lambda_{1},\cdots,\lambda_{n}$ with domains $\Lambda_{1},\cdots,\Lambda_{n}$, the hyperparameter space $\Lambda$ is the cross-product of these domains: $\Lambda=\Lambda_{1}\times\cdots\times\Lambda_{n}$. Given such a hyperparameter space $\Lambda$, we define the hyperparameter optimization problem as follows:
@@ 3.2 Adaptive Arm Filtering
Gaussian UCB takes the statistical data of historical results into consideration. Note that each arm is evaluated multiple times in a single evaluation round. Let NN denotes the number of tries. The formal definition of Gaussian UCB can be expressed as follows:
@@ 3.3 Adaptive Resource Allocation
13:    update res($a$) by subtracting the resource cost in the evaluation process
@@ 5.2 Model Selection
Dataset RR=2 hours, rr=3 RR=4 hours, rr=3
@@ 5.3 Hyperparameter Optimization
BOHB is the state-of-the-art hyperparameter optimization method that outperforms random search [4] and HyperBand [19]. We compared these three methods on two widely-used models: random forest and logistic regression. Similar to model selection, different time limits (22 h and 44 h) are used to evaluate hyperparameter optimization. The hyperparameters in BOASF and BOHB are $c=2,r=3$ and $b_{min}=\frac{1}{27},b_{max}=1,\eta=3$ respectively. The random forest model is configured with 5 hyperparameters:
@@
  • max features, uniform float in $[0.5,1]$.
  • min samples split, uniform integer in $[2,21]$.
  • min samples leaf, uniform integer in $[1,21]$.
@@
  • C, uniform float in $1\mathrm{e}{-4},1\mathrm{e}{4}]$.
  • max iteration, uniform integer in $[50,500]$.
We divided the range of each hyperparameter into 22 intervals. Therefore, BOASF has $2^{5}$ and $2^{3}$ arms respectively. Table 2 and Table 3 shows that BOASF achieved better balanced accuracy score than Hyperopt-full and BOHB on both the logistic regression and random forest settings statistically. The experimental results between BOASF and Hyperopt-full demonstrate that BOASF along with the sub-space partitioning strategy is effective for hyperparameter optimization.
@@ 5.4 Internal Framework Study
Dataset 22 hours 44 hours
@@
Dataset 22 hours 44 hours
@@ 7 DISCUSSION AND CONCLUSION
There are several improvements to do in our method in the future. Firstly, the UCB score works for a random variable range from 0 to 11. Those metrics which do not meet this condition need to be readjusted to this range. Secondly, we use MinMaxScaler to transfer the UCB score to filtering probability since it guarantees a property that the lowest score will be discarded and the highest score will advance to the next round. Other rescaling methods which has this property still work in our method. Thirdly and finally, th TPE method is used to tune the hyperparameters of each arm. We can compare it with other Bayesian Optimizaiton methods.